OpenStudy (anonymous):

a wall can be built by 6 men working 8 hours a day in 5 days. How many days will it take 4 men to build the wall if they wrok only 5 hours a day

OpenStudy (anonymous):

best idea is to figure out the rate per man and then multiply

OpenStudy (anonymous):

more explanation pleease???

OpenStudy (anonymous):

8 hours a day in five day is \(8\times 5=40\) hours it takes 6 men 40 hours to do the job, therefore it takes each man \(6\times 40=120\) hours to do the job

OpenStudy (anonymous):

and so each man's rate is \(\frac{1}{120}\) that is each man does one one hundred and twentieth of a job per hour

OpenStudy (anonymous):

could u calculate the answer i have to make sure u r right...

OpenStudy (anonymous):

now we want the number of days it takes 4 men working 5 hours a day. lets call the number of days it takes \(x\) then \[4\times 5\times \frac{1}{120}\times x=1\] and solve for \(x\)

OpenStudy (anonymous):

ur answer is wrong

OpenStudy (anonymous):

the answer is 6 aaaaand its wrong

OpenStudy (anonymous):

i think 0 is the answer as the wall is already built

OpenStudy (anonymous):

please dont joke...

OpenStudy (anonymous):

iam not joking take problem in physical context matter occupies space and has mass so you cannot make the wall again by just destroying it your question should have been a similar type of wall......or maybe the same dimensions of the wall................

OpenStudy (anonymous):

this is math ......THIS IS PROPORTION...i think ur in the wrong study bro..

OpenStudy (anonymous):

even if it is maths but u cannot exceed your practical limits...............

OpenStudy (anonymous):

if u arent goin to answer leave it to someone else....

OpenStudy (anonymous):

you got angry bro ......please dont mind man..........

OpenStudy (anonymous):

leave the question now i need the help of another person...

OpenStudy (anonymous):

ok tell me whats the answer.. plz

OpenStudy (anonymous):

i have the answer but dunno the method....... it is 12 days

OpenStudy (anonymous):

leave now!

OpenStudy (anonymous):

LEAVE U USELESS PEOPLE I NEED HELP U R DOIN NOTHIING

OpenStudy (anonymous):

the procedure is mentioned below time taken by 6men = 8*5=40hrs work done by 6men= 1/40hrs total work done = 1/(40*6)=1/240 (i) let the 4men take xdays total time taken by four men=5xhrs work done = 1/5x total work done =1/20x (ii) hence from (i) and (ii) 1/20x=1/240 20x=240 x=12 hence they take 12 days

OpenStudy (anonymous):

you get angry very early have patience ..............and above all be happpy...........

OpenStudy (anonymous):

ho ho that is because i wrote \(6\times 40=120\) but in fact \(6\times 40=240\)

OpenStudy (anonymous):

he he..............

OpenStudy (anonymous):

\[4\times 5\times \frac{1}{240}\times x=1\] \[x=\frac{240}{20}=12\] maybe that is the answer

OpenStudy (anonymous):

ok, i am leaving now as requested

OpenStudy (anonymous):

thats what i have solved above methodologically